A monopolist faces the demand curve Q = 60-P/2. The cost function is C = Q^2. Find the output that maximises this monopolist's profits. What are the prices at profits and that output? Find the elasticity of demand at the profit maximising output.
1
Expert's answer
2020-10-09T07:29:46-0400

Q = 60 - "\frac{P}{2}" or P = 120 - 2Q.


Monopolist profit-maximizing quantity is produced, when MR = MC.


"MR = TR' = (P\u00d7Q)' = (120Q - 2Q^{2})' = 120 - 4Q,"

"MC = C' = (Q^{2})' = 2Q."

120 - 4Q = 2Q,


Q = 20 units.

P = 120 - 2"\u00d7" 20 = 80.


"Ed = -b×P/Q = -0.5×\frac{80} {20} = -2,"

so the demand is elastic.